OpenStudy (anonymous):

Help! I'm confused! I'm trying to get m from the equation below but I don't know if I should use that formula or the y-intercept with the x-intercept to find m. I know I'm supposed to use the y-intercept to find the point slope form... D: Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form.

hero (hero):

y = mx + b Insert points into the equation to form a system of equations, then solve for m: 4 = 3m + b -8 = -3m + b

OpenStudy (anonymous):

D: I only know how to do m= y2-y1/x2-x1

hero (hero):

Isolate b in each equation: b = 4 - 3m b = 3m - 8 set b = b: 4 - 3m = 3m -8 Place like terms on the same side, then solve for m: 8 + 4 = 6m 12 = 6m 12/6 = m 2 = m

OpenStudy (anonymous):

oh! okay ;n; I get it now, thanks that help a lot!! :DD

hero (hero):

\[\frac{y_2 - y_1}{x_2 - x_1} = \frac{4 + 8}{3 + 3} = \frac{12}{6} = 2\]

OpenStudy (anonymous):

thaaaaaaaank you so much! :DD

hero (hero):

You only get it when you're able to do these on your own without help. Until then, keep practicing.

OpenStudy (anonymous):

c: will do~